OpenStudy (anonymous):

1 pts. A store sells a 5-pound bag of apples for $4.60. Which of the following is NOT an equivalent price per pound of apples? A. A 20- pound bag for $18.40 B. A 12-pound bag for $11.04 C. A 7-pound bag for $6.44 D. A 15-pound bag for $9.20

OpenStudy (anonymous):

who ever answers this, can you please explain too

OpenStudy (anonymous):

Multiple each amount of pounds by 4.60 and see if its the answer, for example in letter A multiply 20 and 4.6 to see if its = 18.40

OpenStudy (anonymous):

im leaning towards D

OpenStudy (anonymous):

B and C are both not equal.... @RyleeC

OpenStudy (anonymous):

Okay so you now know its now B and C, check for A and D

OpenStudy (anonymous):

not* B and C

OpenStudy (anonymous):

No, it has to be B or C... i need to find the NOT Equavilant

OpenStudy (anonymous):

A and D are equal

OpenStudy (anonymous):

5 pounds = 4.6$ we read it wrong

OpenStudy (anonymous):

Answer is D, 15/5=3*4.6=13.8

OpenStudy (anonymous):

is that right @AnswerMyQuestions

OpenStudy (anonymous):

Ok, since you eliminated B and C, let's check A and D.

OpenStudy (anonymous):

Since you need to multiply 5 x 4 to get 20, you multiply 4.60 x 4 to see if you get 18.40.

OpenStudy (anonymous):

no i didnt eliminate b and c. when i said "both wrong" i meant they both arent equavilant

OpenStudy (anonymous):

What does 4.60 x 4 get you?

OpenStudy (anonymous):

18.4

OpenStudy (anonymous):

so A is wrong. now b c and d

OpenStudy (anonymous):

Yup, and since you multiply 5 x 3 to get 15, see if you get 9.20 when you multiply 4.60 x 3. What do you get?

OpenStudy (anonymous):

ohhh ok so it is D?

OpenStudy (anonymous):

Multiply 4.60 x 3.

OpenStudy (anonymous):

No, not D.

OpenStudy (anonymous):

but its 13.8

OpenStudy (anonymous):

wouldnt that mean it is d?

OpenStudy (anonymous):

Choice D says a A 15-pound bag would be $9.20. But since you multiply 5 x 3 to get 15, you have to multiply 4.60 to 3, which equals 13.8. So a 15-pound bag would actually be $13.8. So D is wrong.

OpenStudy (anonymous):

So D is right..... because we need to find NOT equivilent

OpenStudy (anonymous):

Oh... Yeah, thanks for correcting me.

OpenStudy (anonymous):

phew.... you scared me for a sec lol.... thanks for da help

OpenStudy (anonymous):

Np.